birr 6500 were divided equally among certain number of persons.had there been 15more persons,each would have got birr30less.find the original number of persons.
Answer by macston(5194)

.
x=original number of people
6500/x=(6500/(x+15))+30
6500=6500x/(x+15)+30x
6500(x+15)=6500x+30x(x+15)
6500x+97500=6500x+30x^2+450x
0=30x^2+450x-97500
0=x^2+15x-3250
0=(x-50)(x+65)
x-50=0 OR x+65=0
x=50 OR x=-65
Cannot have -65 people, so the original number of people was 50.
.ANSWER: There were originally 50 people.
.
CHECK:
6500/50=130 birr
6500/(x+15)=6500/(50+15)=6500/65=100 birr
130 birr- 100 birr=30 birr Fifteen more people get 30 birr less.